OpenStudy (love_to_love_you):

What is f(-3) if f(x) = |2+1| +5 a) 0 b) 10 c) 12 d) -10

OpenStudy (love_to_love_you):

Is it d?

OpenStudy (anonymous):

@jigglypuff314 can help you.

jigglypuff314 (jigglypuff314):

Hello @love_to_love_you are you sure you have the right question? I can't seem to find a x in the f(x) = |2+1| + 5 equation

OpenStudy (love_to_love_you):

*2x sorry!

jigglypuff314 (jigglypuff314):

that's better, thank you :) now f(-3) means that you plug in -3 for x so you can get f(-3) = |2(-3)| + 5 from there... = | -6 | + 5 and remember that anything in an absolute values sign | | comes out positive so you can get 6 + 5 = ?

jigglypuff314 (jigglypuff314):

oh wait hmmm... I don't see it listed as any of the given options...

OpenStudy (love_to_love_you):

I don't know why it isn't there.

jigglypuff314 (jigglypuff314):

well could you check the original questions again to see if you made any typos? :/

OpenStudy (anonymous):

it would be d.

jigglypuff314 (jigglypuff314):

@ranbowcrazzzyyy why? :3

jigglypuff314 (jigglypuff314):

ohhhhhhhh I see then |2(-3) + 1| + 5 = |-6+1| + 5 = |-5| + 5 = 5 + 5 = so yes you were correct @love_to_love_you

OpenStudy (love_to_love_you):

Oh, now that makes sense. Well thanks @jigglypuff314 would you mind helping me with a different one?

OpenStudy (love_to_love_you):

@ranbowcrazzzyyy thanks to you too!

OpenStudy (anonymous):

yeah np

OpenStudy (love_to_love_you):

a

OpenStudy (love_to_love_you):

b

OpenStudy (love_to_love_you):

c

OpenStudy (love_to_love_you):

d

jigglypuff314 (jigglypuff314):

slope = rise / run as in how much it goes up over how much it goes to the right slope = 1 = 1/1 so everytime it goes right 1 it goes up 1 slope = 2 = 2/1 so every time it goes right 1 it goes up 2

OpenStudy (anonymous):

ok so @love_to_love_you what els do u need help with?

OpenStudy (love_to_love_you):

so which graph is it @jigglypuff314

jigglypuff314 (jigglypuff314):

slope = 2 = 2/1 so every time it goes right 1 it goes up 2 so which graph has it that everytime you move over 1 box, it goes up 2 boxes?

OpenStudy (love_to_love_you):

b?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (love_to_love_you):

yay! haha thank you!

OpenStudy (love_to_love_you):

So @jigglypuff314 I am doing a study guide, and there are 60 questions and I am stuck on about 20 of them. How many would you be willing to help me with? Hahaha

jigglypuff314 (jigglypuff314):

lol xD I'll try to help you with as many as I can, but you'll have to post each question separately please

OpenStudy (love_to_love_you):

Alright, so close this question, open another, then close that, etc?

OpenStudy (love_to_love_you):

Or stay on this discussion and comment separately? sorry I am new to this.

jigglypuff314 (jigglypuff314):

the first one close open close open close open it's just so we don't spam rainbow with notifications ;)

OpenStudy (love_to_love_you):

ahaha okay!